OpenStudy (anonymous):

can someone help me with this problem? -4+5x=6x+1

OpenStudy (anonymous):

x=-5

OpenStudy (anonymous):

x=-4

OpenStudy (anonymous):

How can x =-4 ?

OpenStudy (anonymous):

can you explain how you got x=-5

OpenStudy (anonymous):

Do you understand equation ?

OpenStudy (anonymous):

@Fool For Math : I Went On My Algebra ..

OpenStudy (anonymous):

no i dont

OpenStudy (anonymous):

then you need to learn it first.

jhonyy9 (jhonyy9):

-4+5X-5x=6X-5x+1 first step you need subtract from both sides 5x and will get -4=x+1 so now you need subtract from both sides 1 and will get -4-1=x+1-1 -5=x - hope so much that this is understandably